OpenStudy (anonymous):

Solve by factoring the trinomial as a binomial square and then use the square root method. Round to the nearest tenth if necessary. x2 + 16x + 64 = 81 A. {9, -9} B. {-8, 9} C. {-17, 1} D. no real solution

OpenStudy (mathteacher1729):

We can use the fact that the "Difference of squares" (google it) :) is \(x^2-a^2=(x+a)(x-a)\) ? In this case \(a^2=64 \Rightarrow a = 8\).

OpenStudy (anonymous):

\[(x+8)^2=81\] \[x+8=9\text { or } x+8=-9\] \[x=1 , \text { or } x = -17\]

OpenStudy (anonymous):

factor \[x^2+16x+64\] as \[(x+8)(x+8)=(x+8)^2\] because \[64=8^2\text { and } 16=2\times 8\]